The product of two even integers is 2020.Find their sum.
1
Expert's answer
2020-07-13T19:13:03-0400

Solution:

2n⋅2m=2020 (n, m in Z+).

n⋅m=505 ⟹ m=505/n⟹ n=1 or n=5 or n=101 or n=505.

n1=1,m1=505; n2=5,m2=101; n3=101,m3=5; n4=505,m4=1⟹ 

sum1=2n1+2m1=1012,

sum2=2n2+2m2=212,

sum3=2n3+2m3=212,

sum4=2n4+2m4=1012.

Answer: sum=1012 or sum=212